PartnerObserving Log for 2007 Feb 13/14

Observer: Tom Herbst, Giacomo Beccari
Inst. Support: John Hill
TOs: Ed Lundin & David Gonzalez

Serious winter storm all afternoon. Accumulation 15-20 cm. We are cooked.

19:00 local time - still snowing

20:30 - snow has stopped...now freezing fog

22:30 - actuator crew has finished working in the mirror cell - installing pin hole test in LBC Blue filter wheel

23:15 - starting pinhole test (to identify excess background) - freezing fog continues...horizontal icicles are now 15-20 cm long

00:00 Happy Valentine's Day - don't forget!

02:15 - snowing big time. Interesting results from pinhole test

03:20 - snow has stopped...now freezing fog again.

-- TomHerbst - 13 Feb 2007
I Attachment Action Size Date Who Comment
092718.jpgjpg 092718.jpg manage 743 K 14 Feb 2007 - 10:32 TomHerbst Pinhole images of scattered light in LBC-Blue, left=B right=V
Todays_Balcony_Pic.jpgjpg Todays_Balcony_Pic.jpg manage 256 K 14 Feb 2007 - 10:27 TomHerbst Todays balcony picture...
Topic revision: r7 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback